OpenStudy (doodleduck8):

The tax on a new car costing $14,400 is $784. At the same tax rate, what is the tax on a vehicle costing $20,000? Round to the nearest dollar. $989 $1,089 $1,568 $1,278 I got $1,089 is that correct if it is can you explain how you got the answer. @amorfide :)

OpenStudy (tkhunny):

That's not how it works. YOU demonstrate how YOU managed that response.

OpenStudy (doodleduck8):

784/14,400 = x/20,000 (20000 * 784) = 14400x x = 20000/14400 = x/784 x = (20000 * 784)/14400 x = $1089 The answer is $1,089

OpenStudy (amorfide):

we know there is a tax rate, so we will have a percentage of 14,400, is equal to 784 so we need to find a way to write this. when you take a percentage of something, let's say 80% of 10 we would change 80% to a decimal, so 80% = 0.8 then we will multiply 10 by this decimal 10*0.8=8 so now that we know that. We know that the value 14,400 is multiplied by something to give 784, we dont know what this something is, so we will let it be a letter, x so 14,400x=784 solve for x

OpenStudy (doodleduck8):

Thanks